# Estimating Heterogeneous Treatment Effects by Combining Weak Instruments and Observational Data

### v1

@@ Abstract
Accurately predicting conditional average treatment effects (CATEs) is crucial in personalized medicine and digital platform analytics. Since often the treatments of interest cannot be directly randomized, observational data is leveraged to learn CATEs, but this approach can incur significant bias from unobserved confounding. One strategy to overcome these limitations is to seek latent quasi-experiments in instrumental variables (IVs) for the treatment, for example, a randomized intent to treat or a randomized product recommendation. This approach, on the other hand, can suffer from low compliance, i.e., IV weakness. Some subgroups may even exhibit zero compliance meaning we cannot instrument for their CATEs at all. In this paper we develop a novel approach to combine IV and observational data to enable reliable CATE estimation in the presence of unobserved confounding in the observational data and low compliance in the IV data, including no compliance for some subgroups. We propose a two-stage framework that first learns biased CATEs from the observational data, and then applies a compliance-weighted correction using IV data, effectively leveraging IV strength variability across covariates. We characterize the convergence rates of our method and validate its effectiveness through a simulation study. Additionally, we demonstrate its utility with real data by analyzing the heterogeneous effects of 401(k) plan participation on wealth.
@@ 1 Introduction
The use of observational data to perform individual-level causal analyses is increasingly commonplace across personalized medicine and online platforms and whenever understanding individualized responses is crucial and/or there is an opportunity to personalize. The key quantity for such analyses is the conditional average treatment effects (CATEs), given baseline covariates (features) with respect to which we want to understand heterogeneity and/or personalize.
@@
Using observational data can nonetheless introduce bias from unobserved confounding, where the observed dependence between outcomes and interventions is not due solely to treatment effects but also to variables that influence both outcome and treatment and are not controlled for by baseline covariates, such as socioeconomic status, healthfulness, busyness, etc. These biases can skew causal effect estimates, resulting in unreliable analysis or even harmful decisions.
@@
Randomized trials are the gold standard for causal inference but are often infeasible. For example, digital services cannot compel a user to see/try/buy a product and clinical trials may not force invasive treatments. An alternative is randomizing the encouragement toward certain actions, such as recommending a product/treatment. These randomized encouragements can serve as instrumental variables (IVs). Under certain conditions, IVs can enable unbiased treatment effect estimation [4].
@@
Identification of CATEs using on IVs crucially hinges on the premise that compliance – the correlation between the treatment received and the intent/encouragement – is nonzero across all baseline-covariate values. And, whenever compliance is nonzero but small, estimation based on IVs exhibits high variance and can become unreliable [3]. In practice, the assumption of strong compliance is often violated. For example, certain users on digital platforms may disregard recommendations either altogether or of certain undesired content, and on mobile health platforms certain participants may ignore recommendations (e.g. taking 250 steps per hour) due to time constraints or disinterest.
@@
To address the challenge of estimating unbiased CATEs in the presence of unobserved confounding and low compliance, we introduce a two-stage framework. In the first stage, we identify a confounded CATE from observational data. Then, in the second stage, we utilize an IV to learn the confounding bias by weighting the samples according to compliance. Assuming only that the bias can be extrapolated, this method extends treatment effect adjustments to those minimally influenced by the IV, employing a transfer learning approach that leverages varying instrument strengths across covariate groups.
@@
This framework mirrors strategies in causal inference that combine randomized trials with observational data to manage low covariate overlap. Building on this body of work, we introduce two methodologies for extrapolating confounding bias within the observational dataset: a parametric estimation approach, assuming the confounding bias adheres to a parametric form, and a transfer learning strategy that assumes a shared representation between the true and biased CATE. We study the properties of our CATE estimators in finite samples and validate our approaches through comprehensive empirical studies.
@@ 2 Related Work
Heterogeneous treatment effect estimation from observational data: Recent advances in machine learning have expanded the use of observational data to estimate CATEs using diverse techniques such as random forests [46], Bayesian algorithms [21], deep learning [43], and meta-learners [30]. However, these methods often unrealistically assume an absence of confounding, limiting their real-world applicability.
@@
Combining observational and randomized data: Increasing research focuses on merging observational datasets with randomized control trial (RCT) data to reduce observational bias. Strategies range from imposing structural assumptions such as strong parametric assumptions [26] or a shared structure between the biased and unbiased CATE functions [20], to optimizing dual estimators from both data types for improved bias correction [50]. Our work aligns with efforts to debias treatment effects using both observational and experimental data, but addresses additional challenges such as low compliance, the need to debias the overall effect function rather than individual outcome functions, and the complexity of estimating CATE from IV data using a ratio estimator.
Where our work lies: To the best of our knowledge, no current estimation technique effectively combines an IV study, particularly one with weak instruments or low compliance, with an observational study to derive robust and unbiased conditional average treatment effects (CATE) estimates. We bridge this gap by introducing two robust and consistent CATE estimation techniques, building upon previous work on combining RCT and observational data [26, 20], as well as work that addresses the complexities associated with weak instruments [11, 1].
@@ 3 Background and Setup
We consider the standard setting of causal inference where we aim to learn the conditional average treatment effect of a binary treatment $A\in\{0,1\}$ on an outcome $Y\in\mathbb{R}$ in the presence of covariates $X\in\mathcal{X}\subseteq\mathbb{R}^{m}$. Our approach is grounded in Rubin’s potential outcomes framework, wherein each unit is associated with two a potential outcomes $Y(0),Y(1)$ of which only $Y=Y(A)$ is observed (causal consistency). Our objective is to learn the CATE function, which is given by:
@@
However, we only have access to $n_{O}$ i.i.d. samples from an observational dataset $O=(X^{O}_{i},A^{O}_{i},Y^{O}_{i})_{i=1}^{n_{O}}\sim(X^{O},A^{O},Y^{O})$. Thus, we are faced with the fundamental problem of causal inference in that we do not observe both potential outcomes, but only the outcome under the administered treatment. Without further assumptions, there exists the possibility of unobserved confounding, leading to a situation where
$\tau^{O}(x)=\mathbb{E}[Y^{O}\mid A^{O}=1,X^{O}=x]-\mathbb{E}[Y\mid A^{O}=0,X^{O}=x]\neq\tau(x),$ (2)
indicating a persistent bias in the observed treatment effects that remains unaffected by increasing the data sample size. We denote this bias by $b(x)$. i.e.:
@@
Alongside the observational dataset, we have $n_{E}$ i.i.d. samples from an experimental, intent-to-treat dataset $E=(X^{E}_{i},Z^{E}_{i},A^{E}_{i},Y^{E}_{i})_{i=1}^{n_{E}}\sim(X^{E},Z^{E},A^{E})$ with a binary instrument $Z^{E}\in\{0,1\}$. We also let $X^{E}\in\mathcal{X}$ and $p_{X^{E}}(x)=p_{X^{O}}(x)$, where $p_{X}$ denotes the density of the random variable $X$. Similar to before, use use $Y^{E}(A,Z)$ to denote the potential outcome with treatment $A$ and instrument $Z$. Furthermore, let $A^{E}(Z)$ denote the potential treatment under instrument $Z$. We assume that this dataset obeys the following standard IV assumptions on the data generating process:
@@
We note that the relevance assumption in Assumption 1 is a weaker version of the standard IV assumptions since we allow for arbitrarily weak instruments in some regions of the covariate spaces. With Assumption 1 and Assumption 2, we can identify the CATE as:
$\tau^{E}(x)=\frac{\mathbb{E}[Y^{E}\mid Z^{E}=1,X^{E}=x]-\mathbb{E}[Y^{E}\mid Z^{E}=0,X^{E}=x]}{\mathbb{E}[A^{E}\mid Z^{E}=1,X^{E}=x]-\mathbb{E}[A^{E}\mid Z^{E}=0,X^{E}=x]}:=\frac{\delta_{Y}(x)}{\gamma(x)}=\tau(x)$ (3)
for $x\in\mathcal{X}^{\prime}$. Here, $\gamma(x)$ denotes heterogeneous compliance, a measure of instrument strength, given by as $\gamma(x)=P(C=1\mid X^{E}=x)$ under Assumption 2. A strong instrument ($\gamma(x)\rightarrow 1$) indicates high adherence to the recommended treatment, with $\gamma(x)=1$ signifying perfect compliance, similar to a true randomized controlled trial (RCT). Conversely, a weak instrument ($\gamma(x)\rightarrow 0$) suggests minimal influence on treatment uptake, with $\gamma(x)=0$ indicating no compliance and a confounded selection into treatment. The relevance assumption in Assumption 1 ensures $\gamma(x)>0$ for $x^{\prime}\in\mathcal{X}^{\prime}$, validating the estimation procedure in Equation 3. However, small $\gamma(x)$ values lead to estimates of $\tau(x)$ with high asymptotic variance. Moreover, we wish to extend $\tau(X)$ estimation from $\mathcal{X}^{\prime}$ to $\mathcal{X}$, our population of interest.
Thus, relying solely on observational data results in biased $\tau(x)$ estimates, while experimental data alone can yield high variance or invalid estimates for $\tau(x)$ with low compliance. This work seeks to overcome these challenges by strategically combining the strengths of both datasets to provide a robust CATE estimation technique.
@@ 4 Estimation Method
To obtain robust CATE estimates for the population $\mathcal{X}$, we propose a two-step framework that integrates information from both the observational data and the IV study. First, we learn the confounded CATE function $\widehat{\tau}^{O}(x)$ using the observational data $(X^{O}_{i},A^{O}_{i},Y^{O}_{i})_{i=1}^{n_{O}}$. This is a common estimation problem in both causal inference and machine learning and can be addressed with many existing techniques including meta-learners ([30]), random forests ([46]), and neural networks ([43]).
Next, we wish to utilize the learned $\widehat{\tau}^{O}(x)$ to approximate the bias function $b(x)=\tau(x)-\tau^{O}(x)$. Without oracle access to the true CATE function $\tau(x)$, we instead use samples from the experimental (IV) study $(X^{E}_{i},A^{E}_{i},Y^{E}_{i})_{i=1}^{n_{E}}$ for which we can estimate an unbiased, but potentially high variance CATE on the compliers. Our data combination strategy hinges on the following lemma:
@@
We note that in the case of randomized instrument assignment, the instrument propensity is known and often given by a constant, i.e. $\pi_{Z}(x)=\pi_{Z}>0$. Letting $V_{Z}(x)=\pi_{Z}(x)(1-\pi_{Z}(x))$, Lemma 1 indicates that we can learn $\gamma(x)$ and, if necessary, $\pi_{Z}(x)$ from data and utilize the pseudo-outcome
@@
in subsequent regression tasks to obtain an unbiased estimate of $\tau(x)$ for $x\in\mathcal{X}^{\prime}$ (provided $\pi_{Z}$ and $\gamma$ are estimated consistently). However, such an estimator requires precise knowledge of the population $\mathcal{X}^{\prime}$ where $\gamma(x)>0$. Additionally, for small values of $\gamma(x)$, $\pi_{Z}(x)$, and $1-\pi_{Z}(x)$, this method may result in high variance in the estimates of $\widehat{b}(x)$, especially for certain parametric function classes. To address these challenges, we weight the data samples by the inverse variance of $\widetilde{Y}^{E}/(\widehat{\gamma}(x)\widehat{V}_{Z}(x))$ given by $\operatorname*{{\rm Var}}(\widetilde{Y}^{E}\lvert X^{E}=x)^{-1}\widehat{\gamma}^{2}(x)\widehat{V}_{Z}^{2}(x)$. This approach is frequently used in generalized least squares methods (GLS, [2]) to confer the algorithm asymptotic efficiency. While $\operatorname*{{\rm Var}}(\widetilde{Y}^{E}\lvert X^{E}=x)$ can be estimated from data using machine learning methods, it is generally preferable to weigh the estimator solely by compliance and instrument propensity to avoid issues with large values of $\operatorname*{{\rm Var}}(\widetilde{Y}^{E}\lvert X^{E}=x)^{-1}$. Assuming the bias function belongs to a class of functions $\mathcal{B}$, our proposed algorithm can be described by the following weighted empirical risk minimization (ERM) procedure.
@@
where the factor $\widehat{\gamma}^{2}(x)\widehat{V}_{Z}^{2}(x)$ was used for weighting the squared loss. This estimator automatically eliminates the need to know $\mathcal{X}^{\prime}$ since we assign weights of 00 when $\widehat{\gamma}=0$. Moreover, this method places higher emphasis on lower-variance pseudo-outcomes, thereby minimizing the risk of overfitting to data points with high variance. This technique is commonly employed in other IV estimation tasks, such as local average treatment effect estimation (LATE), where weighting data points by compliance yields estimators with lower variance ([11, 1]).
The weighting scheme in Equation 4 creates a weighted distribution, $\tilde{p}_{X^{E}}(x)$, for optimizing the ERM procedure. Since $\tilde{p}_{X^{E}}(x)$ differs from the target distribution $p_{X^{E}}(x)$, this creates a transfer learning challenge. Without additional constraints on the function class $\mathcal{B}$, the minimization in Equation 4 may yield many possible solutions. To ensure a unique or limited solution set, $\mathcal{B}$ must have low complexity or require further structural assumptions. We explore two function classes within $\mathcal{B}$: a parametric class defined by $b(x)=\theta^{T}\phi(x)$ with a known mapping $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d}$, and another parametric class where $b(x)=\nu^{T}\phi(x)$ uses $\phi\in\Phi$, a learned representation common to both the observational and IV datasets.

### v2

@@ Abstract
Accurately predicting conditional average treatment effects (CATEs) is crucial in personalized medicine and digital platform analytics. Since the treatments of interest often cannot be directly randomized, observational data is leveraged to learn CATEs, but this approach can incur significant bias from unobserved confounding. One strategy to overcome these limitations is to leverage instrumental variables (IVs) as latent quasi-experiments, such as randomized intent-to-treat assignments or randomized product recommendations. This approach, on the other hand, can suffer from low compliance, i.e., IV weakness. Some subgroups may even exhibit zero compliance, meaning we cannot instrument for their CATEs at all. In this paper, we develop a novel approach to combine IV and observational data to enable reliable CATE estimation in the presence of unobserved confounding in the observational data and low compliance in the IV data, including no compliance for some subgroups. We propose a two-stage framework that first learns biased CATEs from the observational data, and then applies a compliance-weighted correction using IV data, effectively leveraging IV strength variability across covariates. We characterize the convergence rates of our method and validate its effectiveness through a simulation study. Additionally, we demonstrate its utility with real data by analyzing the heterogeneous effects of 401(k) plan participation on wealth.
@@ 1 Introduction
The use of observational data for individual-level causal analyses is becoming increasingly common in personalized medicine, online platforms, and any setting where understanding individualized responses is crucial and/or presents an opportunity for personalization. The key quantity for such analyses is the conditional average treatment effect (CATE), which captures how treatment effects vary according to baseline covariates (features). This measure provides insight into effect heterogeneity and enables personalization.
@@
Using observational data can nonetheless introduce bias from unobserved confounding, where the observed relationship between outcomes and interventions is influenced not only by treatment effects but also by variables that influence both outcome and treatment, such as socioeconomic status, health, user mood, etc., which are not captured by baseline covariates. These biases can skew causal effect estimates, resulting in unreliable analyses or even harmful policy decisions.
@@
Randomized trials are the gold standard for causal inference, but they are often infeasible. For instance, digital services cannot force users to view or buy a product, and clinical trials cannot require invasive treatments. A common alternative is to randomize the encouragement of certain actions, such as recommending a product or treatment. These encouragements can serve as instrumental variables (IVs) which, under certain conditions, enable unbiased estimation of treatment effects [4].
@@
Identification of CATEs using IVs crucially hinges on the premise that compliance – the correlation between the treatment received and the intent/encouragement – is nonzero across all baseline-covariate values. When compliance is nonzero but small, IV-based estimates tend to exhibit high variance, making them unreliable [3]. In practice, the assumption of strong compliance is often violated. For example, users on digital platforms may ignore recommendations entirely or reject certain types of content, while participants on mobile health platforms may disregard prompts (e.g. taking 250 steps per hour) due to time constraints or lack of interest.
@@
To address the challenge of estimating unbiased CATEs in the presence of unobserved confounding and low IV compliance, we introduce a two-stage framework. In the first stage, we estimate a biased, confounded CATE from observational data. Then, in the second stage, we utilize an IV to learn the confounding bias by weighting the samples according to their compliance levels. By assuming only that the bias can be extrapolated, this approach extends treatment effect adjustments even to groups minimally influenced by the IV, employing a transfer learning approach that leverages varying instrument strengths across covariate groups.
@@
This framework mirrors strategies in causal inference that combine randomized trials with observational data to address low covariate overlap. Building on this body of work, we introduce two methodologies for extrapolating confounding bias within the observational dataset: a parametric estimation approach, assuming the confounding bias adheres to a parametric form, and a transfer learning strategy that assumes a shared representation between the true and biased CATE. We study the properties of our CATE estimators in finite samples and validate our approaches through comprehensive empirical studies.
@@ 2 Related Work
Heterogeneous treatment effect estimation from observational data: Recent advances in machine learning have expanded the use of observational data to estimate CATEs using diverse techniques such as random forests [51], Bayesian algorithms [24], deep learning [48], and meta-learners [33]. However, these methods often unrealistically assume an absence of confounding, limiting their real-world applicability. Efforts to account for unobserved confounding either construct bounds on treatment effects [40, 17] or use latent variable models and multiple/sequential treatments to debias CATE estimates [36, 53, 9], but they frequently depend on unverifiable assumptions or require accurate proxy data, reducing their practical utility.
@@
Combining observational and randomized data: Increasing research focuses on integrating observational datasets with randomized control trial (RCT) data to mitigate observational bias. Strategies include imposing structural assumptions, such as strong parametric constraints [29], or assuming a shared structure between biased and unbiased CATE functions [23], as well as optimizing dual estimators from both data types for improved bias correction [55]. Our work aligns with efforts to debias treatment effects using both observational and experimental data, but also addresses challenges such as low IV compliance, the need to debias the overall effect function rather than individual outcome functions, and the complexity of estimating CATEs from IV data using a ratio estimator.
Where our work lies: To the best of our knowledge, no current estimation technique effectively combines an IV study, particularly one with weak instruments or low compliance, with an observational study to derive robust and unbiased CATE estimates. We bridge this gap by introducing two robust and consistent CATE estimation techniques, building upon previous work on combining RCT and observational data [29, 23], as well as work that addresses the complexities associated with weak instruments [13, 1].
@@ 3 Background and Setup
We consider the standard setting of causal inference where the goal is to estimate the conditional average treatment effect of a binary treatment $A\in\{0,1\}$ on an outcome $Y\in\mathbb{R}$ in the presence of covariates $X\in\mathcal{X}\subseteq\mathbb{R}^{m}$. Our approach is grounded in Rubin’s potential outcomes framework, wherein each unit is associated with two potential outcomes $Y(0),Y(1)$ of which only $Y=Y(A)$ is observed (causal consistency). Our objective is to learn the CATE function, which is given by:
@@
However, we only have access to $n_{O}$ i.i.d. samples from an observational dataset $O=(X^{O}_{i},A^{O}_{i},Y^{O}_{i})_{i=1}^{n_{O}}\sim(X^{O},A^{O},Y^{O})$. Thus, we face the fundamental problem of causal inference: only the outcome under the administered treatment is observed, while the counterfactual remains unobserved. Without further assumptions, there exists the possibility of unobserved confounding, leading to a situation where
$\tau^{O}(x)=\mathbb{E}[Y^{O}\mid A^{O}=1,X^{O}=x]-\mathbb{E}[Y^{O}\mid A^{O}=0,X^{O}=x]\neq\tau(x),$ (2)
which indicates a persistent bias in the observed treatment effects that does not diminish even with an increasing sample size. We denote this bias by $b(x)$, that is:
@@
Alongside the observational dataset, we have $n_{E}$ i.i.d. samples from an experimental, intent-to-treat dataset $E=(X^{E}_{i},Z^{E}_{i},A^{E}_{i},Y^{E}_{i})_{i=1}^{n_{E}}\sim(X^{E},Z^{E},A^{E})$ where $Z^{E}$ is a binary instrument taking values in $\{0,1\}$. We let $X^{E}\in\mathcal{X}$ and assume the $p_{X^{E}}(x)=p_{X^{O}}(x)$, where $p_{X}$ denotes the density of the random variable $X$. Moreover, we assume that the joint distribution of covariates and unobserved confounders $(X,U)$ is consistent across both datasets. As before, we use $Y^{E}(A,Z)$ to denote the potential outcome given treatment $A$ and instrument $Z$. Additionally, let $A^{E}(Z)$ denote the potential treatment under instrument $Z$, and define the compliance and defiance indicators $C$ and $D$ by $C:=\mathbb{I}[A^{E}(1)>A^{E}(0)]$ and $D:=\mathbb{I}[A^{E}(1)<A^{E}(0)]$, respectively. We assume that this dataset follows standard IV assumptions on the data generating process:
@@
We note that the relevance assumption in Assumption 1 is a weaker version of the standard IV assumptions since we allow for arbitrarily weak instruments in some regions of the covariate spaces. With Assumption 1 and Assumption 2, we can identify the CATE for $x\in\mathcal{X}^{\prime}$ as:
$\tau^{E}(x)=\frac{\mathbb{E}[Y^{E}\mid Z^{E}=1,X^{E}=x]-\mathbb{E}[Y^{E}\mid Z^{E}=0,X^{E}=x]}{\mathbb{E}[A^{E}\mid Z^{E}=1,X^{E}=x]-\mathbb{E}[A^{E}\mid Z^{E}=0,X^{E}=x]}:=\frac{\delta_{Y}(x)}{\gamma(x)}=\tau(x).$ (3)
We provide the proof of Equation 3 in Appendix B. Here, $\gamma(x)$ denotes heterogeneous compliance, a measure of instrument strength, given by $\gamma(x)=P(C=1\mid X^{E}=x)-P(D=1\mid X^{E}=x)$ under Assumption 2. A strong instrument ($\gamma(x)\rightarrow 1$) indicates high adherence to the recommended treatment, with $\gamma(x)=1$ signifying perfect compliance, similar to a true randomized controlled trial. Conversely, a weak instrument ($\gamma(x)\rightarrow 0$) suggests minimal influence on treatment uptake, with $\gamma(x)=0$ indicating no compliance and a confounded selection into treatment. The relevance assumption in Assumption 1 ensures $\gamma(x)\neq 0$ for $x^{\prime}\in\mathcal{X}^{\prime}$, validating the estimation procedure in Equation 3. However, small $\gamma(x)$ values lead to estimates of $\tau(x)$ with high asymptotic variance. Moreover, we wish to extend the $\tau(x)$ estimation from $\mathcal{X}^{\prime}$ to $\mathcal{X}$, our population of interest.
Thus, relying solely on observational data results in biased $\tau(x)$ estimates, while experimental data alone can yield high variance or invalid estimates for $x\in\mathcal{X}$ with low compliance. This work addresses these challenges by strategically combining the strengths of both datasets to provide a robust CATE estimation technique.
@@ 4 Estimation Method
Figure 1: Illustration of our two-stage procedure: the first stage learns a biased CATE from observational data, while the second stage uses IV data to correct the bias.
To obtain robust estimates of the CATE function for the population of interest $\mathcal{X}$, we propose a two-step framework that integrates information from both the observational data and the IV study. First, we estimate the confounded CATE function $\widehat{\tau}^{O}(x)$ using the observational data $(X^{O}_{i},A^{O}_{i},Y^{O}_{i})_{i=1}^{n_{O}}$. This is a well-established problem in both causal inference and machine learning, and it can be addressed using various existing techniques, including meta-learners ([33]), random forests ([51]), and neural networks ([48]).
Next, we wish to approximate the bias function $b(x)=\tau(x)-\tau^{O}(x)$ using the learned $\widehat{\tau}^{O}(x)$. Without oracle access to the true CATE function $\tau(x)$, we instead rely on samples from the experimental (IV) study $(X^{E}_{i},Z^{E}_{i},A^{E}_{i},Y^{E}_{i})_{i=1}^{n_{E}}$ for which we can estimate an unbiased, though potentially high variance, CATE for $x\in\mathcal{X}^{\prime}$, as given in Equation 3. Our approach hinges on the following lemma:
@@
We note that in the case of randomized instrument assignment, the instrument propensity is known and often given by a constant, i.e., $\pi_{Z}(x)=\pi_{Z}>0$. By defining $V_{Z}(x):=\pi_{Z}(x)(1-\pi_{Z}(x))$, Lemma 1 shows that the bias function $b(x)$ can be expressed in terms of observable quantities as $b(x)=\mathbb{E}\left[\frac{Y^{E}Z^{E}(1-{\pi}_{Z}(X^{E}))-Y^{E}(1-Z^{E}){\pi}_{Z}(X^{E})}{{V}_{Z}(X^{E}){\gamma}(X^{E})}-\tau^{O}(x)\mid X^{E}=x\right]$ for $x\in\mathcal{X}^{\prime}$. This formulation suggests that we can estimate $\gamma(x)$ and, if necessary, $\pi_{Z}(x)$ from data and utilize the pseudo-outcome
@@
along with the estimated $\widehat{\tau}^{O}(x)$, in a subsequent regression task to obtain an unbiased and consistent estimate of $b(x)$ for $x\in\mathcal{X}^{\prime}$ (provided $\pi_{Z}$, $\gamma$, and $\widehat{\tau}^{O}$ are estimated consistently). However, such an estimator only provides estimates for $\mathcal{X}^{\prime}$ where $\gamma(x)\neq 0$. Additionally, for small values of $\gamma(x)$, $\pi_{Z}(x)$, and $1-\pi_{Z}(x)$, this method may result in high variance in the estimates $\widehat{b}(x)$, especially for certain parametric function classes. To address these challenges, we weight the data samples by the inverse variance of $\widetilde{Y}^{E}/(\widehat{\gamma}(x)\widehat{V}_{Z}(x))$ given by $\operatorname*{{\rm Var}}(\widetilde{Y}^{E}\lvert X^{E}=x)^{-1}\widehat{\gamma}^{2}(x)\widehat{V}_{Z}^{2}(x)$. This approach is frequently used in generalized least squares methods (GLS, [2]) to confer the algorithm asymptotic efficiency. While $\operatorname*{{\rm Var}}(\widetilde{Y}^{E}\lvert X^{E}=x)$ can be estimated from data using machine learning methods, it is generally preferable to weight the estimator solely by compliance and instrument propensity to avoid issues with small values of $\operatorname*{{\rm Var}}(\widetilde{Y}^{E}\lvert X^{E}=x)$. Assuming the bias function belongs to a class of functions $\mathcal{B}$, our proposed algorithm can be described by the following weighted empirical risk minimization (ERM) procedure.
@@
where the factor $\widehat{\gamma}^{2}(x)\widehat{V}_{Z}^{2}(x)$ was used for weighting the squared loss. This estimator automatically extrapolates to all of $\mathcal{X}$ since we assign weights of 00 when $\widehat{\gamma}(x)=0$. Moreover, this method places higher emphasis on lower-variance pseudo-outcomes, thereby minimizing the risk of overfitting to data points with high variance. This weighting technique is commonly employed in other IV estimation tasks, such as local average treatment effect estimation (LATE), where weighting data points by compliance yields estimators with lower variance ([13, 1]).
The weighting scheme in Equation 4 creates a weighted distribution, $\tilde{p}_{X^{E}}(x)$, for optimizing the ERM procedure. Since $\tilde{p}_{X^{E}}(x)$ differs from the target distribution $p_{X^{E}}(x)$, this introduces a transfer learning problem. Without additional constraints on the function class $\mathcal{B}$, the minimization in Equation 4 may yield many possible solutions. To ensure a unique or limited solution set, $\mathcal{B}$ must have low complexity or require further structural assumptions. We explore two function classes $\mathcal{B}$: a parametric class defined by $b(x)=\theta^{T}\phi(x),\theta\in\mathbb{R}^{d}$ with a known mapping $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d}$, and a second parametric class where $b(x)=\nu^{T}\phi(x)$, with $\nu\in\mathbb{R}^{d}$ and $\phi\in\Phi$ being a learned representation common to both the observational and IV datasets.